One side of triangle is 3 more than 3 times the shortest side. The third side is 24 feet more than the shortest side. The perimeter is 87. Find all three sides.
Answer by aquacean(7)


a = shortest side
3 + 3a = side b
24 + a = side c
87 = a + 3 + 3a + 24 + a
87 = 5a + 27
87 - 27 = 5a
60 = 5a
60 / 5 = a
a = 12
b = 3 + 3(12)
b = 39
c = 24 + 12
c = 36